OpenStudy (anonymous):

which pair of ratios foms a poportion? A)10/16and5/32 B)10/22and5/11 C)14/8and7/16 D)8/44and16/22

OpenStudy (anonymous):

Equal Ratios form proportions

OpenStudy (anonymous):

To be equal \[\frac{ A}{B } = \frac{ C }{ D }\] one has to have A*D = B*C

OpenStudy (anonymous):

im still a littl confused